Bryan Carpenter A6-5 Slocum Heights Syracuse, NY 13210 USA

June 28, 1996

107 Gainsford Road Bitterne Southampton, SO19 7AW

Dear Mrs Young,

Thank you for inviting me to Justine and David’s wedding. I will be very happy to attend.

Yours sincerely,